Repeated random samples of size 36 are drawn from a population (X) with a mean
µ = 10 and a standard deviation σ = 3. Each sample is used to compute a sample
mean X bar = ∑X / n .
(a) What is the mean of these sample means, i.e., what is the value of µ xbar ?
(b) What is the standard deviation of these sample means, i.e., what is the value of
σ xbar ?
(c) What is the shape of the distribution of these sample means? Why?

Answer by rothauserc(4718)

(a) The mean of the sample means is the mean of the population which is 10
:
(b) The standard deviation of the samples = 3/square root(36) = 3/6 = 1/2
:
(c) Normal distribution since all the samples are greater than 30
: